A 50 kg person launches themselves at a 65 kg box at a speed of 5.8 m/s. The force of friction between the box and the ground is 382 N. How far will the box move?
1
Expert's answer
2019-04-30T09:47:01-0400

First find the total mass m of the person and the box

"m={{m}_{p}}+{{m}_{b}}=50\text{ }kg+65\text{ }kg=115\,kg"

Then we use the Newton's second law of motion to find the deceleration a due to the force F of friction between the box and the ground

"a=-\frac{F}{m}\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,(1)"

The minus sign indicates that they decelerate.

If we know the initial, vi, and final, vf,  speeds of the person and the box, as well as their deceleration, then to find the distance s, that they travels to stop, it is best to use the equation

"s=\frac{v_{f}^{2}-v_{i}^{2}}{2a}"

Since vf=0 m/s we get

"s=\frac{-v_{i}^{2}}{2a}"

Now use equation (1)

"s=\frac{mv_{i}^{2}}{2F}"

Substituting the known values, m=115 kg, vi=5.8 m/s, F=382 N, we get

"s=\frac{115\,kg\cdot {{5.8}^{2}}\text{ }{{m}^{2}}/{{s}^{2}}}{2\cdot 382\text{ }N}\approx 5.1\,m"